# Smooth Regularization for Efficient Video Recognition

### v1

@@ Abstract
We propose a smooth regularization technique that instills a strong temporal inductive bias in video recognition models, particularly benefiting lightweight architectures. Our method encourages smoothness in the intermediate-layer embeddings of consecutive frames by modeling their changes as a Gaussian Random Walk (GRW). This penalizes abrupt representational shifts, thereby promoting low-acceleration solutions that better align with the natural temporal coherence inherent in videos. By leveraging this enforced smoothness, lightweight models can more effectively capture complex temporal dynamics. Applied to such models, our technique yields a $3.8\%$$6.4\%$ accuracy improvement on Kinetics-600. Notably, the MoViNets model family trained with our smooth regularization improves the current state-of-the-art by $3.8\%$$6.1\%$ within their respective FLOP constraints, while MobileNetV3 and the MoViNets-Stream family achieve gains of $4.9\%$$6.4\%$ over prior state-of-the-art models with comparable memory footprints. Our code and models are available at https://github.com/gilgoldm/grw-smoothing.
@@ 1 Introduction
Video recognition has rapidly evolved over the past decade, with models becoming increasingly capable of learning sophisticated spatial and temporal representations. However, despite remarkable advancements, many architectures still suffer from overfitting or inefficient use of temporal information fayyaz20213d; huang2018makes; wang2023videomae. In light of these challenges, we introduce a novel smooth regularization approach designed to instill a strong inductive bias specifically tailored to video data. The key insight behind our method is that video content often exhibits continuous motion and gradual changes in appearance, suggesting that representations should vary smoothly over time. By explicitly encouraging this smoothness, we aim to guide neural networks toward more stable and generalizable internal feature representations, ultimately leading to improved performance across a range of video recognition tasks.
@@ 2 Related Work
Early approaches to video action recognition learned spatiotemporal representations with 3D CNNs carreira2017quo; tran2015learning, with subsequent advances such as Two-Stream networks simonyan2014two and SlowFast feichtenhofer2019slowfast improving the balance between spatial semantics and motion modeling. Temporal segment sampling wang2016temporal provided an efficient way to cover long videos with sparse clips. With the advent of transformers, attention-based models extended self-attention to spatiotemporal tokens bertasius2021space; arnab2021vivit. Subsequent work introduced hierarchical and multiscale designs and localized attention to improve efficiency fan2021multiscale; liu2022videoswin, and explored alternative attention mechanisms such as trajectory attention patrick2021keeping, as well as MLP-like backbones zhang2022morphmlp and video-specific ViT adaptations sharir2021image. Despite strong accuracy, pure transformers often incur substantial computational and memory costs for long videos, particularly relative to compact CNN baselines bertasius2021space; arnab2021vivit; liu2022videoswin.
@@
A parallel line of work targets real-time and on-device deployment by reducing compute and parameters. CNN-based designs dominate this space due to depthwise separable convolutions and efficient temporal operators. MoViNets kondratyuk2021movinets introduce a NAS‑designed family of efficient 3D CNNs and a stream buffer for constant‑memory streaming, achieving strong accuracy–efficiency trade‑offs. X3D feichtenhofer2020x3d systematically compounds network width, depth, and temporal resolution to yield compact 3D CNNs. Temporal Shift Module (TSM) lin2019tsm augments 2D CNN backbones with a zero‑parameter, zero‑FLOP temporal exchange, enabling video recognition with image‑classification‑level compute. Further lightweight temporal modules include TEA li2020tea, TDN wang2021tdn, and TAdaConv huang2021tada, which inject temporal cues with modest overhead. Transformer-based models such as MViT fan2021multiscale and VideoSwin-T liu2022videoswin reduce attention cost via multiscale hierarchies or windowed attention, yet they often remain heavier than the most compact CNN baselines at strict mobile budgets. Hybrid architectures combine convolutions and attention to leverage local inductive bias with global modeling. UniFormer li2022uniformer exemplifies this trend by interleaving convolutional blocks and self-attention. Overall, lightweight video recognition in practice is led by CNNs and convolution–attention hybrids (see Table 1 in the Experiments section below).
@@
An early prominent approach to temporal coherence, slow feature analysis (SFA) wiskott2002slow prefers features that evolve as slowly as possible in time by minimizing the expected squared temporal derivative of each feature subject to zero mean, unit variance, and decorrelation constraints. The motivation is that latent factors in natural videos typically change gradually, so maximally slow features capture stable, semantically meaningful structure.
@@
A complementary line of self-supervised work focuses on verifying or predicting the chronological order of frames or clips. Representative approaches include Shuffle & Learn misra2016shuffle, Sorting Sequences (Order Prediction Networks) lee2017unsupervised, and Odd-One-Out networks fernando2017self. These objectives encourage representations to respect temporal structure by reasoning about sequence order rather than enforcing slowness.
@@ 4 Experiments
Datasets. We report our results on Kinetics-600 (K600) carreira2018short and Kinetics-400 (K400) kay2017kinetics. Both datasets consist of 10-second videos of varying resolutions and frame rates, labeled with 600 and 400 action classes.
@@ 4.1 Results
2 MoViNet-A0kondratyuk2021movinets 72.3 2.7 172 1 $\times$ 50
3 MobileNetV3-Showard2019searching 61.3 2.8 224 1 $\times$ 50
4 MobileNetV3-S+TSMlin2019tsm 65.5 2.8 224 1 $\times$ 50
5 X3D-XSfeichtenhofer2020x3d 70.2 3.9 182 1 $\times$ 20
@@
7 MoViNet-A1kondratyuk2021movinets 76.7 6.0 172 1 $\times$ 50
8 X3D-Sfeichtenhofer2020x3d 73.4 7.8 182 1 $\times$ 40
9 MoViNet-A2kondratyuk2021movinets 78.6 10.3 224 1 $\times$ 50
10 MobileNetV3-Lhoward2019searching 68.1 11.0 224 1 $\times$ 50
11 MobileNetV3-L+TSMlin2019tsm 71.4 11.0 224 1 $\times$ 50
@@
13 X3D-Mfeichtenhofer2020x3d 76.9 19.4 256 1 $\times$ 50
14 X3D-XSfeichtenhofer2020x3d 72.3 23.3 182 30 $\times$ 4
@@
16 MoViNet-A3kondratyuk2021movinets 81.8 56.9 256 1 $\times$ 120
17 X3D-Sfeichtenhofer2020x3d 76.4 76.1 182 30 $\times$ 13
18 X3D-Lfeichtenhofer2020x3d 79.1 77.5 356 1 $\times$ 50
19 MoViNet-A4kondratyuk2021movinets 83.5 105 290 1 $\times$ 80
20 UniFormer-Sli2022uniformer 82.8 167 224 4 $\times$ 16
21 X3D-Mfeichtenhofer2020x3d 78.8 186 256 30 $\times$ 16
22 X3D-Lfeichtenhofer2020x3d 80.7 187 356 1 $\times$ 120
23 I3Dcarreira2018short 71.6 216 224 1 $\times$ 250
24 MoViNet-A5kondratyuk2021movinets 84.3 281 320 1 $\times$ 120
25 MViT-B-16×4fan2021multiscale 82.1 353 224 5 $\times$ 16
26 MoViNet-A6kondratyuk2021movinets 84.8 386 320 1 $\times$ 120
27 UniFormer-Bli2022uniformer 84.0 389 224 4 $\times$ 16
28 XViT (8×)bulat2021space 82.5 425 224 3 $\times$ 8
29 XViT (16×)bulat2021space 84.5 850 224 3 $\times$ 16
30 MViT-B-32×3fan2021multiscale 83.4 850 224 5 $\times$ 32
31 MViTv2-B-32×3li2022mvitv2 85.5 1030 224 5 $\times$ 32
@@
MoViNet-A0kondratyuk2021movinets 65.8 2.7
MoViNet-A2kondratyuk2021movinets 75.0 10.3
@@
MoViNet-A3kondratyuk2021movinets 78.2 56.9
X3D-Sfeichtenhofer2020x3d 73.5 76.1
VideoMambaLi2024VideoMamba 76.9 108

### v2

@@ Abstract
We propose a smooth regularization technique that instills a strong temporal inductive bias in video recognition models, particularly benefiting lightweight architectures. Our method encourages smoothness in the intermediate-layer embeddings of consecutive frames by modeling their changes as a Gaussian Random Walk (GRW). This penalizes abrupt representational shifts, thereby promoting low-acceleration solutions that better align with the natural temporal coherence inherent in videos. By leveraging this enforced smoothness, lightweight models can more effectively capture complex temporal dynamics. Applied to such models, our technique yields a $3.8\%$$6.4\%$ accuracy improvement on Kinetics-600. Notably, the MoViNets model family trained with our smooth regularization improves the current state-of-the-art by $3.8\%$$6.1\%$ within their respective FLOP constraints, while MobileNetV3 and the MoViNets-Stream family achieve gains of $4.9\%$$6.4\%$ over prior state-of-the-art models with comparable memory footprints. Our code and models are available at https://github.com/cmusatyalab/grw-smoothing.
@@ 1 Introduction
Video recognition has rapidly evolved over the past decade, with models becoming increasingly capable of learning sophisticated spatial and temporal representations. However, despite remarkable advancements, many architectures still suffer from overfitting or inefficient use of temporal information Fayyaz et al. (2021); Huang et al. (2018); Wang et al. (2023). In light of these challenges, we introduce a novel smooth regularization approach designed to instill a strong inductive bias specifically tailored to video data. The key insight behind our method is that video content often exhibits continuous motion and gradual changes in appearance, suggesting that representations should vary smoothly over time. By explicitly encouraging this smoothness, we aim to guide neural networks toward more stable and generalizable internal feature representations, ultimately leading to improved performance across a range of video recognition tasks.
@@ 2 Related Work
Early approaches to video action recognition learned spatiotemporal representations with 3D CNNs Carreira and Zisserman (2017); Tran et al. (2015), with subsequent advances such as Two-Stream networks Simonyan and Zisserman (2014) and SlowFast Feichtenhofer et al. (2019) improving the balance between spatial semantics and motion modeling. Temporal segment sampling Wang et al. (2016) provided an efficient way to cover long videos with sparse clips. With the advent of transformers, attention-based models extended self-attention to spatiotemporal tokens Bertasius et al. (2021); Arnab et al. (2021). Subsequent work introduced hierarchical and multiscale designs and localized attention to improve efficiency Fan et al. (2021); Liu et al. (2022), and explored alternative attention mechanisms such as trajectory attention Patrick et al. (2021), as well as MLP-like backbones Zhang et al. (2022) and video-specific ViT adaptations Sharir et al. (2021). Despite strong accuracy, pure transformers often incur substantial computational and memory costs for long videos, particularly relative to compact CNN baselines Bertasius et al. (2021); Arnab et al. (2021); Liu et al. (2022).
@@
A parallel line of work targets real-time and on-device deployment by reducing compute and parameters. CNN-based designs dominate this space due to depthwise separable convolutions and efficient temporal operators. MoViNets Kondratyuk et al. (2021) introduce a NAS‑designed family of efficient 3D CNNs and a stream buffer for constant‑memory streaming, achieving strong accuracy–efficiency trade‑offs. X3D Feichtenhofer (2020) systematically compounds network width, depth, and temporal resolution to yield compact 3D CNNs. Temporal Shift Module (TSM) Lin et al. (2019) augments 2D CNN backbones with a zero‑parameter, zero‑FLOP temporal exchange, enabling video recognition with image‑classification‑level compute. Further lightweight temporal modules include TEA Li et al. (2020), TDN Wang et al. (2021), and TAdaConv Huang et al. (2021), which inject temporal cues with modest overhead. Transformer-based models such as MViT Fan et al. (2021) and VideoSwin-T Liu et al. (2022) reduce attention cost via multiscale hierarchies or windowed attention, yet they often remain heavier than the most compact CNN baselines at strict mobile budgets. Hybrid architectures combine convolutions and attention to leverage local inductive bias with global modeling. UniFormer Li et al. (2022a) exemplifies this trend by interleaving convolutional blocks and self-attention. Overall, lightweight video recognition in practice is led by CNNs and convolution–attention hybrids (see Table 1 in the Experiments section below).
@@
An early prominent approach to temporal coherence, slow feature analysis (SFA) Wiskott and Sejnowski (2002) prefers features that evolve as slowly as possible in time by minimizing the expected squared temporal derivative of each feature subject to zero mean, unit variance, and decorrelation constraints. The motivation is that latent factors in natural videos typically change gradually, so maximally slow features capture stable, semantically meaningful structure.
@@
A complementary line of self-supervised work focuses on verifying or predicting the chronological order of frames or clips. Representative approaches include Shuffle & Learn Misra et al. (2016), Sorting Sequences (Order Prediction Networks) Lee et al. (2017), and Odd-One-Out networks Fernando et al. (2017). These objectives encourage representations to respect temporal structure by reasoning about sequence order rather than enforcing slowness.
@@ 4 Experiments
Datasets. We report our results on Kinetics-600 (K600) Carreira et al. (2018) and Kinetics-400 (K400) Kay et al. (2017). Both datasets consist of 10-second videos of varying resolutions and frame rates, labeled with 600 and 400 action classes.
@@ 4.1 Results
2 MoViNet-A0Kondratyuk et al. (2021) 72.3 2.7 172 1 $\times$ 50
3 MobileNetV3-SHoward et al. (2019) 61.3 2.8 224 1 $\times$ 50
4 MobileNetV3-S+TSMLin et al. (2019) 65.5 2.8 224 1 $\times$ 50
5 X3D-XSFeichtenhofer (2020) 70.2 3.9 182 1 $\times$ 20
@@
7 MoViNet-A1Kondratyuk et al. (2021) 76.7 6.0 172 1 $\times$ 50
8 X3D-SFeichtenhofer (2020) 73.4 7.8 182 1 $\times$ 40
9 MoViNet-A2Kondratyuk et al. (2021) 78.6 10.3 224 1 $\times$ 50
10 MobileNetV3-LHoward et al. (2019) 68.1 11.0 224 1 $\times$ 50
11 MobileNetV3-L+TSMLin et al. (2019) 71.4 11.0 224 1 $\times$ 50
@@
13 X3D-MFeichtenhofer (2020) 76.9 19.4 256 1 $\times$ 50
14 X3D-XSFeichtenhofer (2020) 72.3 23.3 182 30 $\times$ 4
@@
16 MoViNet-A3Kondratyuk et al. (2021) 81.8 56.9 256 1 $\times$ 120
17 X3D-SFeichtenhofer (2020) 76.4 76.1 182 30 $\times$ 13
18 X3D-LFeichtenhofer (2020) 79.1 77.5 356 1 $\times$ 50
19 MoViNet-A4Kondratyuk et al. (2021) 83.5 105 290 1 $\times$ 80
20 UniFormer-SLi et al. (2022a) 82.8 167 224 4 $\times$ 16
21 X3D-MFeichtenhofer (2020) 78.8 186 256 30 $\times$ 16
22 X3D-LFeichtenhofer (2020) 80.7 187 356 1 $\times$ 120
23 I3DCarreira et al. (2018) 71.6 216 224 1 $\times$ 250
24 MoViNet-A5Kondratyuk et al. (2021) 84.3 281 320 1 $\times$ 120
25 MViT-B-16×4Fan et al. (2021) 82.1 353 224 5 $\times$ 16
26 MoViNet-A6Kondratyuk et al. (2021) 84.8 386 320 1 $\times$ 120
27 UniFormer-BLi et al. (2022a) 84.0 389 224 4 $\times$ 16
28 XViT (8×)Bulat et al. (2021) 82.5 425 224 3 $\times$ 8
29 XViT (16×)Bulat et al. (2021) 84.5 850 224 3 $\times$ 16
30 MViT-B-32×3Fan et al. (2021) 83.4 850 224 5 $\times$ 32
31 MViTv2-B-32×3Li et al. (2022b) 85.5 1030 224 5 $\times$ 32
@@
MoViNet-A0Kondratyuk et al. (2021) 65.8 2.7
MoViNet-A2Kondratyuk et al. (2021) 75.0 10.3
@@
MoViNet-A3Kondratyuk et al. (2021) 78.2 56.9
X3D-SFeichtenhofer (2020) 73.5 76.1
VideoMambaLi et al. (2024) 76.9 108